OpenStudy (anonymous):

Solve for x. (x - 5)(x - 5) = 0 -25 -5 5 25 Can I have some steps ? or someones help please ill give you a medal

rvc (rvc):

solve it Open the brackets

OpenStudy (anonymous):

So just remove the brackets then ? Divide ?

OpenStudy (skullpatrol):

Do you understand what it means to "solve" an equation?

jimthompson5910 (jim_thompson5910):

Zero product property if A*B = 0, then either A = 0 or B = 0

jimthompson5910 (jim_thompson5910):

(x - 5)(x - 5) = 0 means either x-5 = 0 or x-5 = 0 another example: (x-2)(x+7) = 0 means either x-2 = 0 or x+7=0

rvc (rvc):

an easy way is put all the values one by one and check which one equals to 0

OpenStudy (skullpatrol):

Specifically, you are looking for those values of x that will make the equation a true statement.

OpenStudy (skullpatrol):

What value of x will make x — 5 = 0 into a true statement?

OpenStudy (anonymous):

Thanks Everyone for all your help and time

rvc (rvc):

\(\huge (x-a)(x-b)=0\) \(\huge(x-a)=0~or~(x-b)=0\)

rvc (rvc):

All the best!!!

OpenStudy (anonymous):

Wish i could everyone a medal but i cant sorry guys ..

OpenStudy (skullpatrol):

The same value of x that makes x = 5 a true statement, namely 5 is also what makes x — 5 = 0 a true statement, right.?